OpenStudy (anonymous):

solving equations 5+3(x-4)=2(x+1)

OpenStudy (anonymous):

distribute:\[5+3x-12=2x+2\]combine like terms:\[3x-7=2x+2\]subtract 2 from both sides:\[3x-9=2x\]subtract 3x from both sides:\[-9=x\] your final answer:\[x=-9\]

OpenStudy (anonymous):

thank you